Let X be a discrete random variable with values in N = {1, 2,...}. Prove that X is geometric with parameter p = P(X = 1) if and only if the memoryless property P(X = n + m | X > n) = P(X = m) holds. To show that the memoryless property implies that X is geometric, you need to prove that the p.m.f. of X has to be P(X = k) = p(1 - p)^(k-1). For this, use P(X = k) = P(X = k + 1|X > 1) repeatedly.
Answer #1

Only -f aut 4irn XGomdriep lX-k)-p( 0941/ New Pxntm /x 7n 3tn 16 1*7) PCX=ntm, X u) . w X 7n) 7CX=ntm) : PCR7n) ntm-1 CHD ken1-(1-1) 71-1 - Plxm) XS K G Lomitrie ) 7)1 m h tant Given xt p m k, n1 PXktX7 1 PlX-k)P(X- k1X7a) PCMT1) 4-11 P(K= kt)rh PIX=H (1x)1C-0- (1-x) (X2 -) P(K 1) PX-3)-) P(x = 2) tuN PCX-4) 1(X-9) 5(4-1)4 0sx 4eonstrie p) where p P(A&1 0:1 Ctrdx